OpenStudy (anonymous):

I neeeeeeeeeeedddddd heeeeeeeelllllllllllllllllppppppppppppp!!!!! The measure of two complementary angles are in the ratio 1 : 9. What are the degree measures of the two angles?

OpenStudy (asnaseer):

complementary angles add up to 90 degrees

OpenStudy (asnaseer):

so, if you call one of them x, then the other one must be 90 - x

OpenStudy (asnaseer):

now use this to express the ratio 1:9 and you should be able to solve for x

OpenStudy (anonymous):

Or you could just multiply ninety by the ratio.

OpenStudy (anonymous):

Let one be x, another one be 9x x+9x = 90 degrees. 10x=90 degrees. x= 90/10 = 9 degrees. Another one is 9x = 9x9=81 degrees. 81+9=90 degrees. There's your answer :)

OpenStudy (anonymous):

okay thanks!!! that really helps!!